One day a movie theatre collected $4275 from 675 people. The price of admission for adults is $7 and a child admission is $5.How many adults and how many children were admitted to the theatre that day?
Answer by checkley71(8405)

A+C=675 OR A=675-C
7A+5C=4275
7(675-C)+5C=4275
4725-7C+5C=4275
-2C=4275-4725
-2C=-450
C=-450/-2
C=225 IS THE NUMBER OF CHILDREN'S TICKETS SOLD.
675-225=450 THE NUMBER OF ADULT TICKETS SOLD.
7*450+5*225=4275
3150+1125=4275
4275=4275